# A Survey of Deep Graph Learning under Distribution Shifts: from Graph Out-of-Distribution Generalization to Adaptation

### v1

@@ 4 Training-Time Graph OOD Adaptation
Training-time graph OOD adaptation focuses on addressing the distribution discrepancies between source and target graph data. By mitigating the distribution shifts, models can effectively transfer knowledge from the source domain and adapt to the target distribution. In general, training-time graph OOD adaptation serves three primary objectives in different scenarios.
@@
  • Cross-graph Knowledge Transfer. To facilitate knowledge transfer from well-labeled graphs to those with limited labels, it is essential to effectively manage distribution shifts between these graphs, as distinct graphs often exhibit varied data distributions.
@@
In this section, we discuss existing training-time graph OOD adaptation methods, highlighting the techniques for mitigating distribution shifts. The summary can be found in Table II.
4.1 Model-Centric Approaches
@@
Invariant Representation Learning. Similar to the invariant representation learning adopted in graph OOD generalization, the goal of this technique in training-time graph OOD adaptation is to learn invariant representations across source and target graphs, ensuring robustness to distribution shifts. This technique is frequently employed for domain adaptation under covariate assumption [^1] Within the context of graph OOD adaptation, we commonly use the terms “distribution” and “domain” interchangeably., in which an invariant relationship between latent representations and labels $\mathcal{P}_{S}(\textbf{Y}|\textbf{H})=\mathcal{P}_{T}(\textbf{Y}|\textbf{H})$ is assumed. Inspired by the theoretical generalization bound [117], domain-invariant representation learning methods aim to train a graph encoder $g(\cdot)$ to minimize discrepancies between the induced marginal source distribution $\mathcal{P}_{S}(\textbf{H})$ and the target distribution $\mathcal{P}_{T}(\textbf{H})$, while also identifying a classifier $f(\cdot)$ in the latent space that minimizes empirical source risk. To achieve these two goals, the loss function for domain-invariant representation learning can be formulated as follows:
@@
  • (1) Distribution Distance Minimization. Distribution distance minimization directly employs the distance between marginal distributions as the regularization term in Equation 5. Methods vary in terms of the choice of the distance metric and the specific representations they aim to align. SR-GNN [118] considers central moment discrepancy as regularization and aligns the distribution discrepancies in the final layer of traditional GCN. CDNE [119], GraphAE [12], and GRADE [120] target at minimizing the statistical discrepancies between source and target across all latent layers, with the regularization term as a summation of distribution distances of different layers. In terms of their distance metric, CDNE uses the marginal maximum mean discrepancy and class-conditional marginal maximum mean discrepancy, GraphAE considers the multiple kernel variant of the maximum mean discrepancy as the distance metric, and GRADE defines and utilizes subtree discrepancy. JHGDA [121] relies on a hierarchical pooling module to extract network hierarchies and minimizes statistical discrepancies in hierarchical representations via the exponential form of the marginal and class-conditional maximum mean discrepancy. For non-trainable representations, such as the latent embeddings in SimpleGCN [2], SR-GNN [118] applies an instance weighting technique. In this approach, the learnable weight parameters are optimized using kernel mean matching to mitigate distribution discrepancies. HC-GST [122] uses pseudo-labels, incorporating heterophily ratios and labeling accuracy to align distributions, with learnable weights optimized via kernel mean matching. Most recently, DREAM [123] employs complementary branches (graph-level and subgraph-enhanced) and enforces consistency between them to mitigate biased pseudo-labels and reduce distribution discrepancies. Addtionally, SelMAG [124] designs an optimal-transport-based algorithm by selecting informative source graphs and adapting both embedding and classification spaces, leveraging meta-learning for precise sub-graph and node selection to improve transferability.
@@
  • (2) Adversarial Learning. Adversarial learning approaches align representations by training the graph encoder $g(\cdot)$ to generate embeddings that confuse the domain discriminator $f_{d}(\cdot)$. Correspondingly, the regularization term in Equation 5 is usually formulated as a minimax game between $g(\cdot)$ and $f_{d}(\cdot)$ as:
    $\min_{f_{d}}\max_{g}l(f_{d}(g(\textbf{A},\textbf{X})),\textbf{Y}_{d}),$
@@
    where $Y_{d}$ denotes the domain label, and $l$ can be chosen as a negative distance loss [125], or a domain classification loss [126, 127, 128, 129, 12, 13, 29]. In addition to framing it as a minimax problem, DANE [130] explores the use of two symmetric and adversarial losses for model training, aiming to achieve bi-directional transfer. Typically, adversarial alignment takes place in the final hidden layer, with the exception being GraphAE [12], which aligns representations in all hidden layers. In addition, it is noteworthy that SGDA [13] also takes the label scarcity issue of the source graph into account by employing a weighted self-supervised pseudo-labeling loss. SDA [29] studies the problem under the open-set setting, utilizing an entropy-based separation strategy to categorize target nodes into certain and uncertain groups and specifically aligning nodes from the certain group with adversarial learning technique. Most recently, JDA-GCN [131] introduces a joint adversarial domain adaptive graph convolutional network, leveraging both local and global graph structures through structural graph alignment, improving the model’s ability to capture intricate dependencies within graph data.
@@
  • (3) Disentangled Representation Learning. Disentangled representation learning separates representations into distinct, informative factors. Among these factors, one is designed to capture domain-invariant features that are essential for the target task (e.g classification). This domain-invariant component remains consistent across different domains to preserve classification-related semantic information, thereby facilitating effective adaptation. And the disentangled components can be selectively manipulated to cater to specific requirements of the target task. The loss function for disentangled representation learning can be defined as:
    $\min_{f,g_{s}}\mathbb{E}_{\textbf{A, X, Y}}[l(f(g_{s}(\textbf{A},\textbf{X})),\textbf{Y})]+\min_{g_{s},g_{o}}(l_{reg}+l_{recon}+l_{add}),$
@@
    where $g_{s}$ denotes the graph encoder for acquiring invariant task-related information, $g_{o}$ denotes graph encoder(s) for other components excluding $g_{s}$, $l_{reg}$ represents a regularization term for enhancing the separation between different components, and $l_{recon}$ denotes a reconstruction loss aiming to recover the original graph structure from the concatenated representation, thereby preventing information loss. Additional terms $l_{add}$ are introduced to facilitate the learning of disentangled representations, enabling specific components to exhibit the desired characteristics. In ASN [132], the representation is decomposed into a domain-private part and a domain-invariant classification-related part. A domain adversarial loss is additionally added to facilitate the learning of invariant representations. Analogous to DIVA [133], DGDA [134] assumes that the graph generation process is controlled independently by domain-invariant semantic latent variables, domain latent variables, and random latent variables. To learn representations with desired characteristics, domain classification loss and noise reconstruction loss are considered as the additional losses.
@@
Model Regularization. Instead of focusing on the process of learning aligned representations, some other methods achieve effective knowledge transfer under distribution shifts through model regularization strategy, including both spectral regularization [16, 137] and knowledge distillation [115, 4]. Building on the derived GNN-based generalization bound, You et al. [16] propose SSReg and MFRReg, which regularize the spectral properties of GNN to enhance transferability. They also extend their theoretical results to the semi-supervised setting with the challenging distribution shifts in $\mathcal{P}(\textbf{Y}|\textbf{A},\textbf{X})$. Liu et al. [137] then delve into the inherent generalization ability of GNN models, proposing the removal of propagation layers in the source graph while stacking multiple propagation layers in the target graph. This model design strategy inherently acts as a regularization for the GNN Lipschitz constant and helps to tighten the generalization bound. Both KDGA [115] and KTGNN [4] employ knowledge distillation strategy and regularize the Kullback–Leibler (KL) divergence between the outputs of teacher and student models. In particular, KDGA aims to address the negative augmentation problem by distilling the knowledge of a teacher model trained on augmented graphs into a partially parameter-shared student model on the original graph. KTGNN, on the other hand, considers the semi-supervised node classification problem for VS-Graph, in which vocal nodes are regarded as the source and silent nodes with incomplete features are regarded as the target. The source classifier and target classifier are respectively constructed, and the knowledge of both of the classifiers is distilled into the student transferable classifier through the KL regularization.
TABLE II: A summary of training-time and test-time graph OOD adaptation methods.
@@
RNA [150] graph Single Concept Maintain discriminative ability
@@
SGOOD [152] Graph Multiple Covariate Maintain discriminative ability
4.2 Data-Centric Approaches
@@
In addition to model-centric strategies, data-centric approaches play a crucial role in addressing training-time graph OOD adaptation challenges. These methods focus on leveraging and enhancing graph data itself to improve adaptation performance. By considering the unique characteristics of graphs, these approaches can effectively address the complexities introduced by distribution shifts. Two prominent data-centric techniques are instance weighting and graph data augmentation.
@@
Instance Weighting. Instance Weighting, which assigns different weights for data points, is a commonly used data-centric technique in traditional transfer learning [8]. Similar strategies of weighting graph instances are also adopted in methods for training-time graph OOD adaptation, including edge weighting [138], node weighting [116], and graph weighting [139, 140] for specific tasks or purposes. Borrowing the idea from Adaboost and TrAda, Ye et al. [138] employ the instance weighting technique for the edge sign prediction task. Edge weights are adjusted in each iteration, with reduced weights assigned to misclassified dissimilar source instances to mitigate distribution shifts across graphs. On the other hand, Liu et al. [116] recognize that the distribution shifts between the original data and the augmented data with pseudo-labels may impede the effectiveness of self-training, and thus propose to reweight the augmented node instances based on information gain to mitigate the gap between the original distribution and the shifted distribution. In multi-source transfer settings, the source graphs may not be equally important for predictions on the target graph and some of them may be of poor quality. To address such challenges, both RSS-GCN [140] and NES-TL [139] employ graph weighting techniques to effectively combine the available source graphs. NES-TL [139] proposes the NES index to quantitatively measure the structural similarity between two graphs, and use the NES-based scores as weights to ensemble weak classifiers trained on instances from each source graph and labeled target instances. On the other hand, RSS-GCN [140] utilizes reinforcement learning to select high-quality source graphs, aiming to minimize the distribution divergence between selected source graphs and target graphs. Such sample selection strategy can be considered as a special binary instance weighting.
@@
Graph Data Augmentation. Different from the aforementioned weighting strategy, several works have delved into exploring graph data augmentation strategies to mitigate distribution shifts, mainly focusing on edge adaptation [141, 142, 143]. Dong et al. [141] find that the dataset shift challenge in edge prediction arises from the presence of links in training and the absence of link observations in testing. To address this, they propose FakeEdge, a subgraph-based link prediction framework that intentionally adds or removes the focal links within the subgraph. This approach aims to decouple the dual role of edges, functioning both as components in representation learning and as labels for link prediction. Bi et al. [142] reconsider the domain-level knowledge transfer problem as learning sample-wise knowledge-enhanced posterior distribution. They first learn the similarities of samples from both source and target graphs and build bridges between each sample and its similar samples containing valuable knowledge for prediction. A GNN model is then employed to transfer knowledge across source and target samples on the constructed bridged-graph. More recently, a novel framework called DC-GST [143] has been introduced to bridge the distribution shifts between augmented training instances and test instances in self-training, which incorporates a distribution-shift-aware edge predictor to improve the model’s generalizability of assigning pseudo-labels, along with a carefully designed criterion for pseudo-label selection. Additionally, LTLP [144] enhances graph data augmentation by addressing the long-tailed distribution of common neighbors in link prediction, effectively increasing the common neighbors of tail node pairs through high-quality edge generation and representation refinement.
5 Test-Time Graph OOD Adaptation
@@
Test-time graph OOD adaptation focuses on adjusting a pre-trained model to new target data that becomes available after the initial training phase. Unlike training-time adaptation, which requires both source and target data simultaneously, test-time adaptation adjusts the model using only the target data. This approach, often referred to as source-free adaptation, plays a crucial role in scenarios where access to source data is restricted. For example, in social networks, source data is typically confidential and inaccessible due to privacy protection purposes and data leakage concerns. In this section, we delve into test-time graph OOD adaptation. A summary for related models is provided in Table II.
5.1 Model-Centric Approaches

### v2

@@ IV Training-Time Graph OOD Adaptation
Training-time graph OOD adaptation is a crucial aspect of addressing distribution discrepancies between source and target graph data. These discrepancies can arise due to various factors, including differences in the underlying graph structures, node features, and edge relationships. Effectively mitigating these distribution shifts is essential for enabling the model to transfer knowledge from the source domain and adapt to the target domain. In general, training-time graph OOD adaptation serves three primary objectives in different settings.
@@
  • Cross-graph Knowledge Transfer. Cross-graph knowledge transfer is a critical challenge when transferring knowledge from well-labeled graphs to graphs with limited or no labels. Distinct graphs often exhibit varied data distributions, leading to discrepancies in the representations learned from each graph. This objective ensures that knowledge learned from the source graph can be effectively applied to the target graph.
@@
In this section, we discuss existing training-time graph OOD adaptation methods, highlighting the techniques for mitigating distribution shifts. The summary can be found in Table III.
IV-A Model-Centric Approaches
@@
Invariant Representation Learning. Unlike the invariant representation learning applied in graph OOD generalization, which focuses on generalizing to unseen target distributions without any access to target domain data, invariant representation learning in training-time graph OOD adaptation explicitly considers both source and target domain data during training. This allows the model to align representations across both domains while addressing distribution shifts. This technique is frequently employed for domain adaptation under covariate assumption [^1] Within the context of graph OOD adaptation, we commonly use the terms “distribution” and “domain” interchangeably., in which an invariant relationship between latent representations and labels $\mathcal{P}_{S}(\textbf{Y}|\textbf{H})=\mathcal{P}_{T}(\textbf{Y}|\textbf{H})$ is assumed. Inspired by the theoretical generalization bound [115], domain-invariant representation learning methods aim to train a graph encoder $g(\cdot)$ to minimize discrepancies between the induced marginal source distribution $\mathcal{P}_{S}(\textbf{H})$ and the target distribution $\mathcal{P}_{T}(\textbf{H})$, while also identifying a classifier $f(\cdot)$ in the latent space that minimizes empirical source risk. To achieve these two goals, the loss function for domain-invariant representation learning can be formulated as follows:
@@
  • (1) Distribution Distance Minimization. Distribution distance minimization directly employs the distance between marginal distributions as the regularization term in Equation 5. Different methods vary in their choice of distance metrics and the specific representations they aim to align. SR-GNN [116] considers central moment discrepancy (CMD) as regularization and aligns the distribution discrepancies in the final layer of traditional GCN. CDNE [117], GraphAE [12], and GRADE [118] target at minimizing the statistical discrepancies between source and target across all latent layers, with the regularization term as a summation of distribution distances of different layers. Regarding the choice of distance metric, CDNE utilizes marginal maximum mean discrepancy and class-conditional marginal maximum mean discrepancy. GraphAE adopts the multiple-kernel variant of maximum mean discrepancy, while GRADE defines and applies subtree discrepancy. JHGDA [119] relies on a hierarchical pooling module to extract network hierarchies and minimizes statistical discrepancies in hierarchical representations via the exponential form of the marginal and class-conditional maximum mean discrepancy. For non-trainable representations, such as the latent embeddings in SimpleGCN [2], SR-GNN [116] applies an instance weighting technique. In this approach, the learnable weight parameters are optimized using kernel mean matching to mitigate distribution discrepancies. HC-GST [120] uses pseudo-labels, incorporating heterophily ratios and labeling accuracy to align distributions, with learnable weights optimized via kernel mean matching. Most recently, DREAM [121] employs complementary branches (graph-level and subgraph-enhanced) and enforces consistency between them to mitigate biased pseudo-labels and reduce distribution discrepancies. Addtionally, SelMAG [122] designs an optimal-transport-based algorithm by selecting informative source graphs and adapting both embedding and classification spaces. It further employs meta-learning to refine subgraph and node selection, improving transferability.
@@
  • (2) Adversarial Learning. Adversarial learning approaches align representations by training the graph encoder $g(\cdot)$ to generate embeddings that confuse the domain discriminator $f_{d}(\cdot)$. Here, the domain discriminator is trained to differentiate between source and target domain representations, while the graph encoder tries to produce embeddings that make this distinction difficult. Correspondingly, the regularization term $l_{reg}$ in Equation 5 is usually formulated as a minimax game between $g(\cdot)$ and $f_{d}(\cdot)$ as:
    $\min_{f_{d}}\max_{g}l(f_{d}(g(\textbf{A},\textbf{X})),\textbf{Y}_{d}),$ (6)
@@
    where $\textbf{Y}_{d}$ denotes the domain label, and $l(\cdot)$ can be chosen as a negative distance loss [123], or a domain classification loss [124, 125, 126, 127, 12, 13, 29]. In addition to framing it as a minimax problem, DANE [128] explores the use of two symmetric and adversarial losses for model training, aiming to achieve bi-directional transfer. Typically, adversarial alignment takes place in the final hidden layer, with the exception being GraphAE [12], which aligns representations in all hidden layers. ASN [129] introduces a decomposition of the learned representation into two components: a domain-private component and a domain-invariant component related to classification. A domain adversarial loss is additionally added to facilitate the learning of invariant representations. In addition, it is noteworthy that SGDA [13] also takes the label scarcity issue of the source graph into account by employing a weighted self-supervised pseudo-labeling loss. SDA [29] studies the problem under the open-set setting, utilizing an entropy-based separation strategy to categorize target nodes into certain and uncertain groups and specifically aligning nodes from the certain group with adversarial learning technique. Most recently, JDA-GCN [130] introduces a joint adversarial domain adaptive graph convolutional network, leveraging both local and global graph structures through structural graph alignment, improving the model’s ability to capture intricate dependencies within graph data.
@@
  • (3) Disentangled Representation Learning. In training-time graph OOD adaptation, the focus is on learning representations that not only separate domain-invariant features but also adapt to the specific distribution shifts between source and target domains. Disentangled representation learning separates representations into distinct, informative factors. Among these factors, one is designed to capture domain-invariant features that are essential for the target task (e.g classification). This domain-invariant component remains consistent across different domains to preserve classification-related semantic information, thereby facilitating effective adaptation. And the disentangled components can be selectively manipulated to cater to specific requirements of the target task. The loss function for disentangled representation learning can be defined as:
    $\min_{f,g_{s}}\mathbb{E}_{\textbf{A, X, Y}}[l(f(g_{s}(\textbf{A},\textbf{X})),\textbf{Y})]+\min_{g_{s},g_{o}}(l_{reg}+l_{recon}+l_{add}),$ (7)
@@
    where $g_{s}$ denotes the graph encoder for acquiring invariant task-related information, while $g_{o}$ refers to encoders for other components excluding $g_{s}$. The term $l_{reg}$ represents a regularization term designed to enhance the separation between different components, $l_{recon}$ is the reconstruction loss used to recover the original graph structure from the concatenated representation, preventing information loss, and $l_{add}$ introduces additional terms that facilitate learning disentangled representations, ensuring that specific components exhibit the desired characteristics. Analogous to DIVA [131], DGDA [132] assumes that the graph generation process is controlled independently by domain-invariant semantic latent variables, domain latent variables, and random latent variables. To learn representations with desired characteristics, domain classification loss and noise reconstruction loss are considered as the additional losses.
@@
Model Regularization. Instead of focusing on the process of learning aligned representations, some other methods achieve effective knowledge transfer under distribution shifts through model regularization strategy, including spectral regularization [16, 135] and knowledge distillation [113, 4]. Building on the derived GNN-based generalization bound, You et al. [16] propose SSReg and MFRReg, which regularize the spectral properties of GNN to enhance transferability. They also extend their theoretical results to the semi-supervised setting with the challenging distribution shifts in $\mathcal{P}(\textbf{Y}|\textbf{A},\textbf{X})$. Liu et al. [135] then delve into the inherent generalization ability of GNN models, suggesting the removal of propagation layers in the source graph while stacking multiple propagation layers in the target graph. This design inherently acts as a regularization for the GNN Lipschitz constant, leading to a tighter generalization bound. Both KDGA [113] and KTGNN [4] employ knowledge distillation, enforcing Kullback–Leibler (KL) divergence regularization between the outputs of teacher and student models. KDGA specifically addresses the negative augmentation problem by distilling knowledge from a teacher model trained on augmented graphs into a partially parameter-shared student model trained on the original graph. In contrast, KTGNN focuses on semi-supervised node classification in VS-Graph, where vocal nodes serve as the source and silent nodes with incomplete features act as the target. Separate classifiers are constructed for the source and target domains, and their knowledge is distilled into a transferable student classifier using KL regularization.
TABLE III: A summary of training-time and test-time graph OOD adaptation methods.
@@
IV-B Data-Centric Approaches
@@
Data-centric approaches for training-time graph OOD adaptation leverage the unique structure and properties of graph data to address distribution shifts. These methods enhance the graph’s input data by considering its inherent topological and relational features, which are crucial for effective adaptation across domains. Two prominent techniques in this context are instance weighting and graph data augmentation, both designed to mitigate distribution shifts and improve model performance.
@@
Instance Weighting. Instance Weighting, which assigns different weights to data points, is a commonly used data-centric technique in traditional transfer learning [8]. Similar strategies have been applied to graph-based OOD adaptation, where instance weighting methods include edge weighting [136], node weighting [114], and graph weighting [137, 138], each designed for specific tasks. Inspired by Adaboost and TrAda, Ye et al. [136] employ instance weighting for the edge sign prediction task. They iteratively adjust edge weights, decreasing the weights of misclassified dissimilar source instances to mitigate distribution shifts across graphs. On the other hand, Liu et al. [114] recognize that the distribution shifts between the original data and the augmented data with pseudo-labels may impede the effectiveness of self-training, and thus propose to reweight the augmented node instances based on information gain to mitigate the gap between the original distribution and the shifted distribution.
In multi-source transfer settings, not all source graphs contribute equally to predictions on the target graph, and some may be of low quality. To address this, both RSS-GCN [138] and NES-TL [137] employ graph weighting techniques to effectively integrate multiple source graphs. NES-TL [137] proposes the NES index to quantitatively measure the structural similarity between two graphs, and uses the NES-based scores as weights to ensemble weak classifiers trained on instances from each source graph along with labeled target instances. Additionally, RSS-GCN [138] utilizes reinforcement learning to select high-quality source graphs, aiming to minimize the distribution divergence between selected source and target graphs. This sample selection strategy can be considered as a special case of binary instance weighting.
@@
Graph Data Augmentation. Unlike the instance weighting strategy, which assigns weights to graph components, graph data augmentation methods modify graph structures to mitigate distribution shifts. It is important to clarify that this discussion focuses on graph data augmentation for training-time adaptation rather than generalization, as the former specifically aims to address distribution shifts during training, mainly focusing on edge adaptation. Dong et al. [139] identify that distribution shifts in edge prediction stem from the presence of links in training but their absence during testing. To address this, they propose FakeEdge, a subgraph-based link prediction framework that intentionally adds or removes the focal links within the subgraph. This method separates the dual role of edges, treating them both as features for representation learning and as labels for link prediction. Bi et al. [140] approach the domain-level knowledge transfer problem by redefining it as learning a sample-wise knowledge-enhanced posterior distribution. In their framework, they compute similarities between source and target samples, constructing bridges that connect each sample to its similar samples containing valuable knowledge for prediction. A GNN model is then trained to transfer knowledge across source and target samples using this bridged-graph. This approach allows for smoother adaptation by explicitly bridging domain discrepancies between the source and target graphs, ensuring that knowledge from the source domain is transferred to the target domain effectively. More recently, a novel framework called DC-GST [141] has been introduced to bridge the distribution shifts between augmented training instances and test instances in self-training, which incorporates a distribution-shift-aware edge predictor to improve the model’s generalizability of assigning pseudo-labels, along with a carefully designed criterion for pseudo-label selection. Additionally, LTLP [142] enhances graph data augmentation by addressing the long-tailed distribution of common neighbors in link prediction. It achieves this by generating high-quality edges to increase the number of common neighbors for tail node pairs and refining their representations.
V Test-Time Graph OOD Adaptation
@@
Test-time graph OOD adaptation focuses on adjusting a pre-trained model to new target data that becomes available after the initial training phase. Unlike training-time adaptation, which requires simultaneous access to both source and target data, test-time adaptation updates the model using only the target data. This approach, often referred to as source-free adaptation, is particularly important in scenarios where access to source data is restricted. For instance, in social networks, source data is typically confidential and inaccessible due to privacy protection regulations and concerns about data leakage. In this section, we delve into test-time graph OOD adaptation. A summary for related models is provided in Table III.
V-A Model-Centric Approaches